Consider the titration of 80.0 mL of 0.0200 M C6H5NH2 (a weak base; Kb = 4.30e-10) with 0.100 M HBrO4. Calculate the pH after the following volumes of titrant have been added:


(a) 0.0 mL

pH =  



(b) 4.0 mL

pH =  



(c) 8.0 mL

pH =  



(d) 12.0 mL

pH =  



(e) 16.0 mL

pH =  



(f) 25.6 mL

pH =  
Answer #1


1)when 0.0 mL of HBrO4 is added
C6H5NH2 dissociates as:

C6H5NH2 +H2O     ----->     C6H5NH3+   +   OH-
2*10^-2                   0         0
2*10^-2-x                 x         x


Kb = [C6H5NH3+][OH-]/[C6H5NH2]
Kb = x*x/(c-x)
Assuming x can be ignored as compared to c
So, above expression becomes

Kb = x*x/(c)
so, x = sqrt (Kb*c)
x = sqrt ((4.3*10^-10)*2*10^-2) = 2.933*10^-6

since c is much greater than x, our assumption is correct
so, x = 2.933*10^-6 M



So, [OH-] = x = 2.933*10^-6 M


use:
pOH = -log [OH-]
= -log (2.933*10^-6)
= 5.5328


use:
PH = 14 - pOH
= 14 - 5.5328
= 8.4672
Answer: 8.47

2)when 4.0 mL of HBrO4 is added
Given:
M(HBrO4) = 0.1 M
V(HBrO4) = 4 mL
M(C6H5NH2) = 0.02 M
V(C6H5NH2) = 80 mL


mol(HBrO4) = M(HBrO4) * V(HBrO4)
mol(HBrO4) = 0.1 M * 4 mL = 0.4 mmol

mol(C6H5NH2) = M(C6H5NH2) * V(C6H5NH2)
mol(C6H5NH2) = 0.02 M * 80 mL = 1.6 mmol



We have:
mol(HBrO4) = 0.4 mmol
mol(C6H5NH2) = 1.6 mmol

0.4 mmol of both will react
excess C6H5NH2 remaining = 1.2 mmol
Volume of Solution = 4 + 80 = 84 mL
[C6H5NH2] = 1.2 mmol/84 mL = 0.0143 M
[C6H5NH3+] = 0.4 mmol/84 mL = 0.0048 M

They form basic buffer
base is C6H5NH2
conjugate acid is C6H5NH3+

Kb = 4.3*10^-10

pKb = - log (Kb)
= - log(4.3*10^-10)
= 9.367

use:
pOH = pKb + log {[conjugate acid]/[base]}
= 9.367+ log {4.762*10^-3/1.429*10^-2}
= 8.889

use:
PH = 14 - pOH
= 14 - 8.8894
= 5.1106

Answer: 5.11

3)when 8.0 mL of HBrO4 is added
Given:
M(HBrO4) = 0.1 M
V(HBrO4) = 8 mL
M(C6H5NH2) = 0.02 M
V(C6H5NH2) = 80 mL


mol(HBrO4) = M(HBrO4) * V(HBrO4)
mol(HBrO4) = 0.1 M * 8 mL = 0.8 mmol

mol(C6H5NH2) = M(C6H5NH2) * V(C6H5NH2)
mol(C6H5NH2) = 0.02 M * 80 mL = 1.6 mmol



We have:
mol(HBrO4) = 0.8 mmol
mol(C6H5NH2) = 1.6 mmol

0.8 mmol of both will react
excess C6H5NH2 remaining = 0.8 mmol
Volume of Solution = 8 + 80 = 88 mL
[C6H5NH2] = 0.8 mmol/88 mL = 0.0091 M
[C6H5NH3+] = 0.8 mmol/88 mL = 0.0091 M

They form basic buffer
base is C6H5NH2
conjugate acid is C6H5NH3+

Kb = 4.3*10^-10

pKb = - log (Kb)
= - log(4.3*10^-10)
= 9.367

use:
pOH = pKb + log {[conjugate acid]/[base]}
= 9.367+ log {9.091*10^-3/9.091*10^-3}
= 9.367

use:
PH = 14 - pOH
= 14 - 9.3665
= 4.6335

Answer: 4.63

4)when 12.0 mL of HBrO4 is added
Given:
M(HBrO4) = 0.1 M
V(HBrO4) = 12 mL
M(C6H5NH2) = 0.02 M
V(C6H5NH2) = 80 mL


mol(HBrO4) = M(HBrO4) * V(HBrO4)
mol(HBrO4) = 0.1 M * 12 mL = 1.2 mmol

mol(C6H5NH2) = M(C6H5NH2) * V(C6H5NH2)
mol(C6H5NH2) = 0.02 M * 80 mL = 1.6 mmol



We have:
mol(HBrO4) = 1.2 mmol
mol(C6H5NH2) = 1.6 mmol

1.2 mmol of both will react
excess C6H5NH2 remaining = 0.4 mmol
Volume of Solution = 12 + 80 = 92 mL
[C6H5NH2] = 0.4 mmol/92 mL = 0.0043 M
[C6H5NH3+] = 1.2 mmol/92 mL = 0.013 M

They form basic buffer
base is C6H5NH2
conjugate acid is C6H5NH3+

Kb = 4.3*10^-10

pKb = - log (Kb)
= - log(4.3*10^-10)
= 9.367

use:
pOH = pKb + log {[conjugate acid]/[base]}
= 9.367+ log {1.304*10^-2/4.348*10^-3}
= 9.844

use:
PH = 14 - pOH
= 14 - 9.8437
= 4.1563

Answer: 4.16

5)when 16.0 mL of HBrO4 is added
Given:
M(HBrO4) = 0.1 M
V(HBrO4) = 16 mL
M(C6H5NH2) = 0.02 M
V(C6H5NH2) = 80 mL


mol(HBrO4) = M(HBrO4) * V(HBrO4)
mol(HBrO4) = 0.1 M * 16 mL = 1.6 mmol

mol(C6H5NH2) = M(C6H5NH2) * V(C6H5NH2)
mol(C6H5NH2) = 0.02 M * 80 mL = 1.6 mmol



We have:
mol(HBrO4) = 1.6 mmol
mol(C6H5NH2) = 1.6 mmol

1.6 mmol of both will react to form C6H5NH3+ and H2O
C6H5NH3+ here is acid
C6H5NH3+ formed = 1.6 mmol
Volume of Solution = 16 + 80 = 96 mL

use:
Ka = Kw/Kb
Kw is dissociation constant of water whose value is 1.0*10^-14 at 25 oC
Ka = (1.0*10^-14)/Kb
Ka = (1.0*10^-14)/4.3*10^-10
Ka = 2.326*10^-5
concentration ofC6H5NH3+,c = 1.6 mmol/96 mL = 0.0167 M


C6H5NH3+      + H2O ----->     C6H5NH2   +   H+
1.667*10^-2                    0         0
1.667*10^-2-x                  x         x


Ka = [H+][C6H5NH2]/[C6H5NH3+]
Ka = x*x/(c-x)
Assuming x can be ignored as compared to c
So, above expression becomes

Ka = x*x/(c)
so, x = sqrt (Ka*c)
x = sqrt ((2.326*10^-5)*1.667*10^-2) = 6.226*10^-4

since x is comparable c, our assumption is not correct
we need to solve this using Quadratic equation
Ka = x*x/(c-x)
2.326*10^-5 = x^2/(1.667*10^-2-x)
3.876*10^-7 - 2.326*10^-5 *x = x^2
x^2 + 2.326*10^-5 *x-3.876*10^-7 = 0
This is quadratic equation (ax^2+bx+c=0)
a = 1
b = 2.326*10^-5
c = -3.876*10^-7

Roots can be found by
x = {-b + sqrt(b^2-4*a*c)}/2a
x = {-b - sqrt(b^2-4*a*c)}/2a

b^2-4*a*c = 1.551*10^-6

roots are :
x = 6.111*10^-4 and x = -6.343*10^-4

since x can't be negative, the possible value of x is
x = 6.111*10^-4

[H+] = x = 6.111*10^-4 M

use:
pH = -log [H+]
= -log (6.111*10^-4)
= 3.2139
Answer: 3.21

6)when 25.6 mL of HBrO4 is added
Given:
M(HBrO4) = 0.1 M
V(HBrO4) = 25.6 mL
M(C6H5NH2) = 0.02 M
V(C6H5NH2) = 80 mL


mol(HBrO4) = M(HBrO4) * V(HBrO4)
mol(HBrO4) = 0.1 M * 25.6 mL = 2.56 mmol

mol(C6H5NH2) = M(C6H5NH2) * V(C6H5NH2)
mol(C6H5NH2) = 0.02 M * 80 mL = 1.6 mmol



We have:
mol(HBrO4) = 2.56 mmol
mol(C6H5NH2) = 1.6 mmol

1.6 mmol of both will react
excess HBrO4 remaining = 0.96 mmol
Volume of Solution = 25.6 + 80 = 105.6 mL
[H+] = 0.96 mmol/105.6 mL = 0.0091 M

use:
pH = -log [H+]
= -log (9.091*10^-3)
= 2.0414
Answer: 2.04